OpenStudy (anonymous):

Trig / Pre Cal/ identities Am I on the right path? Please guide me and do not give me the answers \[ (\tan x + \sin x)(1-cos x) = \sin^2 x \tan x \] Do I multiply like with foil starting with tan or do I turn tan into \( \frac{sin x}{cox x} \) and then use foil?

Nnesha (nnesha):

foil! :-) i guess that will be great

OpenStudy (anonymous):

Just with tan \[ \tan x -(\tan x)(\cos x) + \sin x -(\sin x)(\cos x) \] correct so far?

Nnesha (nnesha):

yep looks right :-)

OpenStudy (anonymous):

Ok, how do I approch it from here? do I turn all the tan into sin / cos? and work it out? \[ \tan x -(\tan x)(\cos x) + \sin x -(\sin x)(\cos x) \]

Nnesha (nnesha):

yes convert 2nd tan to sin /cos and let me find notebook and a pencil :3 hehe

Nnesha (nnesha):

make sure question is right :3 :-)

OpenStudy (anonymous):

Yes it is right :-)

Nnesha (nnesha):

is it equal to tanxsinx or tan^2xsinx ??

OpenStudy (anonymous):

I have \[ \tan x -\sin x + \sin x -(\sin x)(\cos x) = \sin^2 x \tan x \] \[ \tan x -(\sin x)(\cos x) = \sin^2 x \tan x \]

Nnesha (nnesha):

yeah so for the final answer i got sinx tanx so that's why ..hm are u sure it's sin square ??

Nnesha (nnesha):

next step is to change tan to sin/cos

OpenStudy (anonymous):

Yes, it is sin^2 x tan x

Nnesha (nnesha):

:( alright

Nnesha (nnesha):

alright got it!! :-)

Nnesha (nnesha):

hahah i'm soo happy :P

OpenStudy (anonymous):

\( \tan x -(\sin x)(\cos x) = \sin^2 x \tan x \) \( \frac{sin x}{cos x} -(\sin x)(\cos x) = \sin^2 x \tan x \) ????

Nnesha (nnesha):

\( \huge\rm \frac{sin x}{cos x} -(\sin x)(\cos x) = \) ???? find common denominator :-)

OpenStudy (anonymous):

Got ya :-)

OpenStudy (anonymous):

one sec

Nnesha (nnesha):

alright :-)

OpenStudy (anonymous):

I am using cos as the common denominator. that is correct?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Ok I got \[ \frac{sin x}{cos x} -(\sin x)(\cos x) = \sin^2 x \tan x \] \[ \frac{(sin x)(cos x)}{\cos x} -\frac{((\sin x))}{\cos x} \frac{\cos^2}{cos} = \sin^2 x \tan x \] correct???

Nnesha (nnesha):

\[ \frac{(sin x)(\cancel{cos x})}{\cos x} -\frac{((\sin x))}{\cos x} \frac{\cos^2}{cos} = \sin^2 x \tan x \] multiply cosx/1 so when you multiply first fraction by cos both cos will cancel each other out

Nnesha (nnesha):

so there isn't supposed to be any cosx \[ \frac{(sin x)}{\cos x} -\frac{((\sin x cos^2x))}{cos x} = \sin^2 x \tan x \] like this

OpenStudy (anonymous):

Oh right!!! Made a boo boo

OpenStudy (anonymous):

One sec, now we should have

OpenStudy (anonymous):

\[ \frac{(sin x)}{\cos x} -\frac{((\sin x cos^2x))}{cos x} = \sin^2 x \tan x \] \[ \frac{(sin x)}{\cos x} -\frac{((\sin x (1-\sin^2 x)2x))}{cos x} = \sin^2 x \tan x \] Ok I am lost

OpenStudy (anonymous):

\[ \frac{(sin x)}{\cos x} -\frac{((\sin x (1-cos^2 x))}{cos x} = \sin^2 x \tan x \]

Nnesha (nnesha):

\[\cos x \times \frac{ \sin x }{ \cos x} - \frac{ sinx \cos x }{ \cos } \times \cos \] multiply